# Language Model Preference Evaluation with Multiple Weak Evaluators

### v2

@@ Abstract
Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge. Existing works usually leverage a powerful LLM (e.g., GPT4) as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is vulnerable to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results. To improve model-based preference evaluation, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results. In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure. We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure. Extensive experiments across ten benchmark datasets show that GED outperforms baseline methods in model ranking, response selection, and model alignment tasks. Notably, GED combines weaker evaluators like Llama3-8B, Mistral-7B, and Qwen2-7B to surpass the performance of stronger evaluators like Qwen2-72B, highlighting its ability to enhance evaluation reliability and improve model performance.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
Figure 1: The figure is divided into two parts. The upper part shows the construction of a preference graph from pairwise LLM evaluations, with cycles indicating inconsistent preferences highlighted in red. The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024). These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, revealing a significant cycle rate (details in Appendix A.5) in their evaluations. This highlights the limitations of even advanced LLMs as reliable evaluators.
@@
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A). These cyclic patterns introduce inconsistencies that undermine the reliability of the evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. The whole process we show in the upper part of Figure 1. Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators (Refer to Appendix A.3 for further details).
@@
To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations. Our method involves two key steps: (1) ensembling multiple weak evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph. By aggregating evaluations from multiple weak evaluators, we ”average out” the noise and biases, resulting in a more robust approximation of the true preference structure. The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks. We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
@@
To validate the practical efficacy of GED, we conduct extensive experiments across model ranking, response selection, and model alignment tasks, utilizing ten widely recognized benchmark datasets, including HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), GAIA (Mialon et al., 2023), LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022). In these experiments, GED consistently outperformed baseline methods. For example, in the response selection task, applying GED yielded an average improvement of 4.51% compared to baseline methods across multiple benchmarks. Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from weaker evaluators surpassed the performance of even stronger individual evaluators. For instance, by using Llama3-8B, Mistral-7B, and Qwen2-7B as weaker evaluators, GED exceeded the performance of using the stronger Qwen2-72B model as an evaluator. These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
@@ 2.1 Preference Graph
A preference graph is defined as a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents a set of $n$ alternatives or candidates, $A\subseteq V\times V$ is a set of directed arcs representing pairwise preferences between these alternatives, and $w:A\rightarrow\mathbb{R}^{+}$ is a weight function that assigns a positive real value to each arc, indicating the strength of the preference.
For any pair of distinct vertices $u,v\in V$, an arc $(u,v)\in A$ exists if there is a preference for $u$ over $v$, with the weight $w(u,v)$ reflecting the intensity of this preference. Formally, this can be represented as:
$(u,v)\in A\quad\text{if and only if}\quad w(u,v)>0$ (1)
The weight function $w(u,v)$ aggregates individual preferences or scores for the pair $(u,v)$. If multiple preference sources exist, the weight can be expressed as:
$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u))$ (2)
where $s_{i}(u,v)$ is the score or preference from the $i$-th source. The preference graph encapsulates the aggregate preferences among all pairs of alternatives, with the weight of each arc representing the cumulative preference strength derived from underlying data or models.
Figure 2: The GED framework mitigates noise and inconsistencies in preference evaluations by utilizing multiple weak evaluators and applying a graph denoising process. It consists of three stages: (1) Graph Ensemble, which combines individual preference graphs; (2) Graph Denoising, which removes cycles to ensure an acyclic structure; and (3) Graph-to-Ranking, which extracts a reliable ranking from the denoised graph.
@@ 2.2 GED: Preference Graph Ensemble and Denoise
As illustrated in Figure 2, our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs. It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking. The detailed steps are as follows:
@@
In the context of graph ensemble, given multiple weighted graphs $G_{1}=(V,A_{1},w_{1}),G_{2}=(V,A_{2},w_{2}),\ldots,G_{k}=(V,A_{k},w_{k})$ that share the same set of vertices $V$ but may differ in their arc sets $A_{i}$, the goal is to combine these graphs into a single ensemble graph $G_{E}=(V,A_{E},w_{E})$. The ensemble graph $G_{E}$ is formed by first defining the arc set $A_{E}$ as the union of the arc sets of the individual graphs and the weight function $w_{E}:A_{E}\rightarrow\mathbb{R}^{+}$ for the ensemble graph is then determined by summing the weights of the corresponding arcs in each graph.
@@
Graph denoising involves transforming the original graph $G=(V,A,w)$ into a DAG. This transformation is achieved by identifying and removing a set of arcs known as the Feedback Arc Set (FAS) (Gabow, 1995), which is a set of arcs whose removal makes the graph acyclic. The goal is to find a minimum FAS, denoted as $R^{*}(G)$, which is a set of arcs with the smallest total weights that needs to be removed to eliminate all cycles in $G$.
To find this minimum FAS, we can order the vertices of $G$ in a specific sequence $s=\{v_{1},v_{2},\ldots,v_{n}\}$. This vertex sequence induces a FAS $R(s)$, consisting of all arcs that point against the direction of the sequence, i.e., arcs $v_{j}\rightarrow v_{i}$ where $j>i$. The graph denoising problem is thus reframed as finding an optimal vertex sequence $s^{*}$ that induces the minimal FAS, such that $R(s^{*})=R^{*}(G)$. This optimal sequence $s^{*}$ ensures that the total weights of arcs eliminated to achieve a DAG is minimized.
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix L.
@@
Given a DAG graph $G=(V,A,w)$, our goal is to derive a ranking based on the structure of $G$. For each vertex $v\in V$, we compute the descendant count $\text{desc}(v)$, defined as the number of vertices that are reachable from $v$ through directed arcs:
$\text{desc}(v)=\left|\{u\in V:v\rightsquigarrow u\}\right|,$ (3)
where $v\rightsquigarrow u$ denotes that there is a directed path from $v$ to $u$. Vertices are then ranked based on their descendant counts, with higher descendant counts indicating a higher position in the ranking. Formally, the ranking $\mathcal{R}$ is represented as a sequence of subsets:
=$\mathcal{R}=\{v_{1},v_{2},\ldots,v_{n}\},$ (4)
where each $v_{i}$ represents a set of vertices with the $i$-th highest descendant count. The final ranking is then:
$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$ (5)
This structured approach ensures that the ranking reflects not only the individual preferences captured in the graph but also the relative strength of these preferences as represented through their descendant connections.
@@ 2.3 Applications
We apply GED to three tasks: Response Selection, selecting the best response from LLM-generated candidates; Model Ranking, ranking models based on task performance; and Model Alignment, identifying the best instruction-response pairs for training. The steps are as follows:
@@
In the response selection task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, with the objective of identifying the optimal answer $ans_{q}^{*}$ for each query. To achieve this, we employ multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$, who provide pairwise preferences among the candidate answers. For each question $q$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the candidate answers, while the directed arcs $A_{a}$ indicate the pairwise preferences among these responses. Each arc is weighted by $w_{a}$, reflecting the strength of the preference indicated by the evaluator. The construction of preference graphs involves evaluating each pair of candidate answers $ans_{i}$ and $ans_{j}$. Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference. This process is detailed in Appendix D. After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs. The aggregation begins with merging all preference graphs into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, which is subsequently processed to remove cycles, resulting in a DAG. From this DAG, we derive the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ of the candidate answers, where the highest-ranked answer is selected as $ans_{q}^{*}$. This process is repeated for each question $q\in Q$, yielding a final set of selected answers $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$. This set reflects a consensus from multiple evaluators in $\mathcal{A}$, ensuring the chosen answers represent the highest quality responses based on rigorous preference evaluation.
@@
In model ranking task, the goal is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to a series of questions $Q=\{q_{1},\ldots,q_{t}\}$. A group of evaluators, denoted as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$, assesses the model outputs by providing preferences for pairs of responses for each question. For each question $q\in Q$, the evaluators generate preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$ over the models. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the models, while the directed arcs $A_{a}$ indicate pairwise preferences, with weights $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$ reflecting the strength of these preferences. The preference graph for a given question $q$ is constructed by evaluating each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, represented by nodes $v_{i}$ and $v_{j}$. For their respective answers $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$, evaluators provide a preference indicating which answer is favored. This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength. The detailed procedure is outlined in Appendix D. Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs. The method begins by merging all preference graphs into a single graph $G_{q}=(V_{q},A_{q},w_{q})$, which is then transformed into a DAG by removing cycles. The final ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ is derived from this DAG. This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix M. This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
@@
In the model alignment task, we have multiple data pairs of the form $(x,y_{1}),(x,y_{2}),\ldots,(x,y_{n})$ for each instruction $x$. The objective is to identify the best response $y^{*}$ corresponding to each instruction $x$. We utilize multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ to provide pairwise preferences among the candidate responses. For each instruction $x$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{x},A_{a},w_{a})$ represents the preferences of evaluator $a$. The vertices $V_{x}$ correspond to the candidate responses $\{y_{1},\ldots,y_{n}\}$, and the directed arcs in $A_{a}$ indicate preferences, weighted by $w_{a}$. After constructing the preference graphs, we apply GED to aggregate them into a single graph $G_{x}$. This graph undergoes a denoising process to remove cycles, allowing us to derive a ranking $\mathcal{R}_{x}$ of the responses. The highest-ranked response in $\mathcal{R}_{x}$ is selected as $y^{*}$ for that instruction $x$. This process is repeated for all $t$ instructions, resulting in the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, which reflects a consensus across the evaluators in $\mathcal{A}$.

### v3

@@ Abstract
Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge. Existing works usually leverage an LLM as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is weak evaluator due to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results. To address this, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results. In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure. We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure. Extensive experiments on ten benchmarks demonstrate GED’s superiority in three applications: model ranking, response selection, and model alignment tasks. Notably, GED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024; Boyeau et al., 2024; Chatzi et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
Figure 1: (a) A preference graph exhibiting cyclic inconsistencies (e.g., A $\succ$ B $\succ$ C $\succ$ A), which violate transitivity. (b) Empirical results showing that even advanced LLMs (e.g., GPT-4-o) exhibit significant noise in preference judgments, leading to inconsistent evaluations. (c) Overview of our proposed framework, GED, which ensembles multiple weak evaluators and applies denoising to recover a directed acyclic graph.
@@
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B, Response B is better than Response C, yet paradoxically also rank Response C as better than Response A. These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al., 2022; Song et al., 2024; Hou et al., 2024; Liu et al., 2024), thereby undermining the reliability of evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. This process is illustrated in Figure 1 (a). Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. In this paper, we define an evaluator as weak if it produces cycles into the preference graph. As shown in Figure  1 (b), we evaluated 10 Llama3-70B (AI@Meta, 2024) responses on HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B (AI@Meta, 2024) as evaluators. Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, highlighting persistent noise and the limitations of these models as weak evaluators.
@@
To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations. Our method involves two key steps: (1) ensembling multiple weak evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph. By aggregating evaluations from multiple weak evaluators, we ”average out” the noise and biases, resulting in a more robust approximation of the true preference structure. The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks. The overall process of GED is illustrated in Figure 1 (c). We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
@@
To validate the practical efficacy of GED, we conduct extensive experiments across model ranking, response selection, and model alignment tasks, utilizing ten widely recognized benchmark datasets, including HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), GAIA (Mialon et al., 2023), LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022). In these experiments, GED consistently outperformed baseline methods. For instance, in the response selection task, GED achieved an average performance gain of 4.51% over baseline methods across multiple benchmarks. Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators. Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, GED exceeded the performance of using the Qwen2-72B in response selection task. These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
2 Related Work
Preference evaluation of LLMs.
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013; Boyeau et al., 2024; Chatzi et al., 2024; Shankar et al., 2024; Naresh et al., 2024), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
Weak supervision.
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b, 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
@@ 3.1 Preference Graph
A preference graph is a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents $n$ candidates, $A\subseteq V\times V$ is a set of directed arcs indicating pairwise preferences, and $w:A\rightarrow\mathbb{R}^{+}$ assigns weights to arcs, representing preference strength.
For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$, where the weight $w(u,v)$ aggregates individual preferences:
$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u)),$
with $s_{i}(u,v)$ representing the score from the $i$-th source. The preference graph encapsulates aggregate preferences, with arc weights reflecting cumulative strength.
@@ 3.2 GED: Preference Graph Ensemble and Denoise
Our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs. It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking. The detailed steps are as follows:
@@
Given $k$ weighted graphs $G_{1},\dots,G_{k}$ with shared vertex set $V$, the ensemble graph $G_{E}=(V,A_{E},w_{E})$ is constructed by setting $A_{E}=\bigcup_{i=1}^{k}A_{i}$ and defining the weight $w_{E}(u,v)=\sum_{i=1}^{k}w_{i}(u,v)$ for each arc $(u,v)\in A_{E}$.
@@
Graph denoising involves transforming the original graph $G=(V,A,w)$ into a DAG. This transformation is achieved by identifying and removing a set of arcs known as the Feedback Arc Set (FAS) (Gabow, 1995), which is a set of arcs whose removal makes the graph acyclic. The goal is to find a minimum FAS, denoted as $R^{*}(G)$, which is a set of arcs with the smallest total weights that needs to be removed to eliminate all cycles in $G$. To find this minimum FAS, we can order the vertices of $G$ in a specific sequence $s=\{v_{1},v_{2},\ldots,v_{n}\}$. This vertex sequence induces a FAS $R(s)$, consisting of all arcs that point against the direction of the sequence, i.e., arcs $v_{j}\rightarrow v_{i}$ where $j>i$. The graph denoising problem is thus reframed as finding an optimal vertex sequence $s^{*}$ that induces the minimal FAS, such that $R(s^{*})=R^{*}(G)$. This optimal sequence $s^{*}$ ensures that the total weights of arcs eliminated to achieve a DAG is minimized.
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix M.
@@
Given a DAG $G=(V,A,w)$, we derive a ranking by computing the descendant count $\text{desc}(v)$ for each vertex $v$, defined as the number of vertices reachable from $v$:
$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$
where $v\rightarrow u$ denotes a directed path. Vertices are ranked based on $\text{desc}(v)$, with ties broken lexicographically:
$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$
This ranking reflects both individual preferences and their relative strengths in the graph.
@@ 3.3 Applications
We apply GED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training).
@@
In this task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, and the goal is to identify the optimal answer $ans_{q}^{*}$. Multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences among these candidates, constructing a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$.
Each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents the preferences of evaluator $a$, where $V_{q}$ corresponds to the candidates and $A_{a}$ indicates pairwise preferences with weights $w_{a}$ reflecting preference strength. This process is detailed in Appendix E. To aggregate these graphs, GED first merges them into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, removes cycles to obtain a DAG, and derives the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$. The top-ranked answer is selected as $ans_{q}^{*}$. Repeating this for all $q\in Q$ yields the final set $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$, representing consensus from multiple evaluators and ensuring high-quality responses.
@@
The goal of model ranking is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to questions $Q=\{q_{1},\ldots,q_{t}\}$. Evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences for model outputs, constructing preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for each question $q$.
Here, each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents preferences, where $V_{q}$ corresponds to the models and $w_{a}$ reflects preference strength. The detailed procedure is outlined in Appendix E. Using GED, we aggregate these graphs into a unified graph $G_{q}$, transform it into a DAG by removing cycles, and derive the ranking $\mathcal{R}_{q}$. Repeating this for all $q\in Q$ yields a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. Finally, a ranking ensemble is applied to compute the overall ranking $\mathcal{R}^{*}$ across questions, reflecting model performance as assessed by evaluators, as detailed in Appendix N
@@
In the model alignment task, the objective is to identify the best response $y^{*}$ for each instruction $x$ from candidate responses $\{y_{1},\ldots,y_{n}\}$. Evaluators $\mathcal{A}$ provide pairwise preferences, forming preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where $G_{a}=(V_{x},A_{a},w_{a})$ represents preferences over $\{y_{1},\ldots,y_{n}\}$. GED aggregates these graphs into a unified graph $G_{x}$, removes cycles to produce a DAG, and derives a ranking $\mathcal{R}_{x}$. The highest-ranked response is selected as $y^{*}$ for $x$. Repeating this for all $t$ instructions yields the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, reflecting consensus among evaluators.